Order of magnitude of (2)30 is approximately as raised power of 10
 

Grade:11

2 Answers

Dheeraj
5 years ago
Let y=2^30
Take log both side
Logy=30*log2
(Log2=0.30)
Logy=30*0.30=9
Take antilog
Y=10^9
So order of magnitude of 2^30 is approx 9 raised power of 10
ISRAR AHAMED HR
2 years ago
Let y = 2^30
Take log On both side
Log y= 30 log 2
Wkt log2=.30
Therefore 10^9
 
30*.30=9
Take anti log